# Mixing Time Bounds for the Gibbs Sampler under Isoperimetry

### v1

@@ 4.1 Three-set isoperimetric inequalities via functional inequalities
$\Upsilon(t)=\frac{\mathtt{s}_{q}t^{q}}{1+\left(1+e^{-1}\right)\mathtt{s}_{q}t^{q}},\qquad\Psi(t)=\frac{t}{2}\cdot\log\left(\frac{2}{t}\right).$
@@
As seen before, the map $x\mapsto\mathcal{D}(x,S_{1})$ is Lipschitz continuous with constant 1. It follows that the function $f$ is equal to 1 on $S_{1}$, vanishes on $S_{2}$, and is Lipschitz continuous on a subset of $S_{3}$, with the modulus of its gradient equal to $\mathcal{D}(S_{1},S_{2})^{-1}$; elsewhere on $S_{3}$, $f$ vanishes. In particular, $f(x)\in[0,1]$ for $x\in S_{3}$. Applying (3) to $f$, we obtain
$\mathbb{E}\left[\left\lvert f\right\rvert^{q}\cdot\log\left\lvert f\right\rvert^{q}\right]-\mathbb{E}\left[\left\lvert f\right\rvert^{q}\right]\cdot\log\mathbb{E}\left[\left\lvert f\right\rvert^{q}\right]\leq\frac{1}{\mathtt{s}_{q}}\cdot\mathbb{E}\left[\left\lvert\nabla f(x)\right\rvert^{q}\right]\leq\frac{\pi(S_{3})}{\mathtt{s}_{q}\cdot\mathcal{D}(S_{1},S_{2})^{q}}.$ (7)
For $x\in[0,1]$, consider the function $\rho(x)=x\cdot\log(1/x)$. Using the convention that $00\cdot\log(0)=1\cdot\log(1)=$ 0, we can rewrite
$\mathbb{E}\left[\left\lvert f\right\rvert^{q}\cdot\log\left\lvert f\right\rvert^{q}\right]={\int_{S_{3}}}\left\lvert f(x)\right\rvert^{q}\cdot\log\left(\left\lvert f(x)\right\rvert^{q}\right)\pi(\mathrm{d}x)=-{\int_{S_{3}}}\rho\left(\left\lvert f(x)\right\rvert^{q}\right)\pi(\mathrm{d}x)$
@@
$\mathbb{E}\left[\left\lvert f\right\rvert^{q}\right]\cdot\log\mathbb{E}\left[\left\lvert f\right\rvert^{q}\right]$ $=\left(\pi(S_{1})+\int_{S_{3}}\left\lvert f(x)\right\rvert^{q}\pi(\mathrm{d}x)\right)\cdot\log\left(\pi(S_{1})+\int_{S_{3}}\left\lvert f(x)\right\rvert^{q}\pi(\mathrm{d}x)\right)$
@@
$\mathbb{E}\left[\left\lvert f\right\rvert^{q}\cdot\log\left\lvert f\right\rvert^{q}\right]\geq-e^{-1}\cdot\pi(S_{3}).$ (8)
@@
$-\mathbb{E}\left[\left\lvert f\right\rvert^{q}\right]\cdot\log\mathbb{E}\left[\left\lvert f\right\rvert^{q}\right]\geq\pi(S_{1})\cdot\log\left(\frac{1}{\pi(S_{1})}\right)-\pi(S_{3}).$ (9)
@@
$\pi(S_{3})$ $\geq\frac{\mathcal{D}(S_{1},S_{2})^{q}}{\mathtt{s}_{q}^{-1}+(1+e^{-1})\cdot\mathcal{D}(S_{1},S_{2})^{q}}\cdot\max\left\{\pi(S_{1})\cdot\log\left(\frac{1}{\pi(S_{1})}\right),~{}\pi(S_{2})\cdot\log\left(\frac{1}{\pi(S_{2})}\right)\right\}$
$\geq\frac{\mathcal{D}(S_{1},S_{2})^{q}}{\mathtt{s}_{q}^{-1}+(1+e^{-1})\cdot\mathcal{D}(S_{1},S_{2})^{q}}\cdot\min\left\{\pi(S_{1}),\pi(S_{2})\right\}\cdot\log\left(\frac{1}{\min\left\{\pi(S_{1}),\pi(S_{2})\right\}}\right)$
$\geq\frac{\mathcal{D}(S_{1},S_{2})^{q}}{\mathtt{s}_{q}^{-1}+(1+e^{-1})\cdot\mathcal{D}(S_{1},S_{2})^{q}}\cdot\frac{\min\left\{\pi(S_{1}),\pi(S_{2})\right\}}{2}\cdot\log\left(\frac{2}{\min\left\{\pi(S_{1}),\pi(S_{2})\right\}}\right)$
@@
Notably, the $L^{q}$-Poincaré and $L^{q}$-log-Sobolev inequalities both lead to three-set isoperimetric inequalities with $\Upsilon(t)\sim\mathtt{p}_{q}t^{q}$ as $t\to 0$. Here the sign $\sim$ means that the two quantities are asymptotic, meaning their ratio converges to a constant as $t\to 0$. This regime is particularly relevant in high-dimensional settings, where $t$ typically scales inversely with the dimension, as per the discussion following Corollary 3.11. In contrast, when $\pi$ is log-concave, we have $\Upsilon(t)\sim t$ as $t\to 0$ (Andrieu et al., 2024, Subsection 3.2).
@@ 4.2 Close coupling conditions for the Gibbs kernels
Here, we establish close coupling conditions for the systematic-scan and random-scan Gibbs kernels, assuming that the conditional distributions of the target distribution $\pi$ are TV continuous (as per Definition 2.5). The following results rely on sequential coupling arguments; see Roberts and Rosenthal (2013) for a formal construction of such couplings, and Marton (2004); Wang and Wu (2014) for related coupling techniques in Wasserstein distance. We begin with the systematic-scan Gibbs sampler.
@@
$\delta=\left(Md\right)^{-1/\beta},\qquad\varepsilon=e^{-1}.$
@@
By construction, we have $U\sim P_{\mathrm{SS}}(x,\cdot)$ and $V\sim P_{\mathrm{SS}}(y,\cdot)$. Moreover, observe that for all $k\in[d]$, we have $U^{\left(k\right)}=\left(U_{1:k},x_{k+1:d}\right)$ and $V^{\left(k\right)}=\left(V_{1:k},y_{k+1:d}\right)$. Using this, (10) can be rewritten as
@@
We now turn our attention towards obtaining a close coupling for the random-scan Gibbs sampler. However, one issue immediately becomes apparent: when $x,y\in\mathbb{R}^{d}$ do not share a coordinate (i.e., when $x-y$ has a non-zero component in each coordinate direction), the supports of $P_{\mathrm{RS}}(x,\cdot)$ and $P_{\mathrm{RS}}(y,\cdot)$ are disjoint. In this case,
@@
[…]
@@
Let $I_{1},\dots,I_{N}\sim\mathrm{Unif}[d]$ be independent and identically distributed. Denote by $\Xi$ the set of unique values attained by $I_{1},\dots,I_{N}$. If $N=\lceil 4d\cdot\log d\rceil$, then
@@
Lemma 4.5 is closely related to the classic coupon collector problem in probability (Erdős and Rényi, 1961). Using the same notation, observe that when $N=d$, we have
@@
as $d\to\infty$, by Stirling’s formula. The probability of observing all $d$ unique values in exactly $d$ draws thus decreases exponentially with the dimension $d$.
@@
Suppose that the conditional distributions of $\pi$ are $(M,\beta)$-TV continuous. Let $N_{d}=\lceil 4d\cdot\log d\rceil$. Then,
$\left\|P_{\mathrm{RS}}^{N_{d}}(x,\cdot)-P_{\mathrm{RS}}^{N_{d}}(y,\cdot)\right\|_{\textsc{TV}}\leq 1-\frac{1}{2}(1-M\left\lvert x-y\right\rvert^{\beta})^{N_{d}}\qquad\forall x,y\in\mathbb{R}^{d}.$
@@
$\delta=\left(MN_{d}\right)^{-1/\beta},\qquad\varepsilon=\left(2e\right)^{-1}.$
@@
By construction, $U\sim P_{\mathrm{RS}}^{N_{d}}(x,\cdot)$ and $V\sim P_{\mathrm{RS}}^{N_{d}}(y,\cdot)$. As before, denote by $\Xi$ the set of unique values attained by $I_{1},\dots,I_{N_{d}}$. When $x$ and $y$ differ in at least two coordinates, the event $\left\{U=V\right\}$ has a non-zero probability of occurring only if $\Xi=[d]$, i.e., when all coordinates have been updated at least once by $P_{\mathrm{RS}}^{N_{d}}$. More generally, for all $x,y\in\mathbb{R}^{d}$, we have
@@
By Lemma 4.5, $\mathbb{P}(\Xi=\left[d\right])\geq 1/2$ when $N_{d}=\left\lceil 4d\cdot\log d\right\rceil$. Hence, for this value of $N_{d}$, we get
@@
$\leq M\left\lvert x-y\right\rvert^{\beta},$ (15)
where the subscript in the expectation indicates that it is taken with respect to the joint distribution of $U^{(k-1)},V^{(k-1)},I_{k}$. Combining (13), (14), (15), we finally obtain
@@ 4.3 Mixing time bounds for the Gibbs sampler
By combining the three-set isoperimetric inequalities from Subsection 4.1 with the close coupling conditions from Subsection 4.2, we can bound the conductance of the systematic-scan and random-scan Gibbs kernels. These results apply to target distributions that satisfy a $L^{q}$-Poincaré or $L^{q}$-log-Sobolev inequality, and whose conditional distributions are TV continuous. Throughout this subsection, we use $\lesssim$ to denote inequalities that hold up to a universal constant factor.
@@
Consider a target distribution $\pi$ that satisfies $\mathrm{PI}_{q}$ with constant $\mathtt{p}_{q}$. Suppose that the conditional distributions of $\pi$ are $(M,\beta)$-TV continuous. Then, the conductance of the systematic-scan Gibbs kernel is bounded by
$\Phi(P_{\mathrm{SS}})\gtrsim\min\left\{1,\frac{\mathtt{p}_{q}}{\left(Md\right)^{q/\beta}}\right\}.$
@@
$\tau(\zeta,P_{\mathrm{SS}})\lesssim\max\left\{1,\frac{\left(Md\right)^{2q/\beta}}{\mathtt{p}_{q}^{2}}\right\}\cdot\log\left(\frac{\sqrt{\Omega}}{\zeta}\right).$
@@
The bound on the conductance follows from combining Corollary 3.11, Proposition 4.1, and Proposition 4.4. The bound on the mixing time is obtained via Corollary 3.5. ∎
@@
Consider a target distribution $\pi$ that satisfies $\mathrm{PI}_{q}$ with constant $\mathtt{p}_{q}$. Suppose that the conditional distributions of $\pi$ are $(M,\beta)$-TV continuous. Let $N_{d}=\left\lceil 4d\cdot\log d\right\rceil$. Then, the conductance of the iterated random-scan Gibbs kernel is bounded by
$\Phi\left(P^{N_{d}}_{\mathrm{RS}}\right)\gtrsim\min\left\{1,\frac{\mathtt{p}_{q}}{\left(MN_{d}\right)^{q/\beta}}\right\}.$
@@
$\lambda_{2}\left(P^{N_{d}}_{\mathrm{RS}}\right)\gtrsim\min\left\{1,\frac{\mathtt{p}_{q}^{2}}{\left(MN_{d}\right)^{2q/\beta}}\right\}.$
@@
$\tau\left(\zeta,P^{N_{d}}_{\mathrm{RS}}\right)\lesssim\max\left\{1,\frac{\left(MN_{d}\right)^{2q/\beta}}{\mathtt{p}_{q}^{2}}\right\}\cdot\log\left(\frac{\sqrt{\Omega}}{\zeta}\right).$
@@
The conductance bound follows from Corollary 3.11, Proposition 4.1, and Proposition 4.6. The spectral gap is then bounded using Theorem 3.7, while the mixing time bound is obtained via Corollary 3.5. ∎
We now focus on bounding the conductance and spectral gap of the random-scan Gibbs kernel $P_{\mathrm{RS}}$. As noted by Narayanan and Srivastava (2022, Lemma 5.3), for any reversible Markov kernel $P$ and for all $N\geq 1$, we have
@@
Let $\sigma(P)$ denote the second largest eigenvalue of $P$, so that $\lambda_{2}(P)=1-\sigma(P)$. Notably, $\sigma\left(P^{N}\right)=\sigma(P)^{N}$, which implies that
@@
$\Phi(P_{\mathrm{RS}})\gtrsim\frac{1}{d\cdot\log d}\cdot\min\left\{1,\frac{\mathtt{p}_{q}}{M^{q/\beta}(d\cdot\log d)^{q/\beta}}\right\}.$
@@
$\lambda_{2}(P_{\mathrm{RS}})\gtrsim\frac{1}{d\cdot\log d}\cdot\min\left\{1,\frac{\mathtt{p}_{q}^{2}}{M^{2q/\beta}(d\cdot\log d)^{2q/\beta}}\right\}.$
@@
$\tau(\zeta,P_{\mathrm{RS}})\lesssim d\cdot\log d\cdot\max\left\{1,\frac{M^{2q/\beta}(d\cdot\log d)^{2q/\beta}}{\mathtt{p}_{q}^{2}}\right\}\cdot\log\left(\frac{\sqrt{\Omega}}{\zeta}\right).$
These results demonstrate that the systematic-scan and random-scan Gibbs samplers exhibit polynomial mixing time dependence on the dimension $d$ for a large class of target distributions. Note that a single step of the systematic-scan kernel $P_{\mathrm{SS}}$ updates all $d$ coordinates, whereas the random-scan kernel $P_{\mathrm{RS}}$ only updates one. Even after accounting for this difference, a residual $\log d$ factor remains in the mixing time bounds for random-scan. This additional $\log d$ factor highlights a modest penalty compared to systematic scan for well-behaved target distributions. This aligns with prior heuristic observations about the efficiency of these sampling schemes (Diaconis, 2013) and arises naturally from spectral bounds on mixing times (Gaitonde and Mossel, 2024).
By Proposition 4.3, all results in this subsection extend to distributions satisfying $\mathrm{LSI}_{q}$ instead of $\mathrm{PI}_{q}$, with the conductance bounds remaining equivalent up to constant factors. Consequently, the bounds on the mixing time in total variation distance remain unchanged when replacing a Poincaré inequality with a log-Sobolev inequality of the same order. However, improvements may be possible when using a different norm to define the mixing time. For instance, in $\chi^{2}$-divergence, Chen et al. (2020, Theorem 1) demonstrated that the logarithmic factor in $\Psi$ from Proposition 4.3 can be exploited to improve the mixing time’s dependence on the warm start parameter (see Corollary 3.5), reducing it from $\log$ to $\log\log$.

### v2

@@ 4.1 Three-set isoperimetric inequalities via functional inequalities
As established above, the function $f$ is Lipschitz continuous with $\left\lvert\nabla f(x)\right\rvert\leq\mathcal{D}(S_{1},S_{2})^{-1}$ for almost every $x\in\mathbb{R}^{d}$. Moreover, $f$ takes the constant value 1 on $S_{1}$, vanishes on $S_{2}$, and satisfies $f(x)\in[0,1]$ for all $x\in S_{3}$. Applying (4) to $f$, we obtain
$\mathbb{E}[\left\lvert f\right\rvert^{q}\cdot\log\left\lvert f\right\rvert^{q}]-\mathbb{E}[\left\lvert f\right\rvert^{q}]\cdot\log\mathbb{E}[\left\lvert f\right\rvert^{q}]\leq\frac{1}{\mathtt{s}_{q}}\cdot\mathbb{E}[\left\lvert\nabla f(x)\right\rvert^{q}]\leq\frac{\pi(S_{3})}{\mathtt{s}_{q}\cdot\mathcal{D}(S_{1},S_{2})^{q}}.$ (9)
For $x\in[0,1]$, consider the function $\rho(x)=x\cdot\log(1/x)$. We can rewrite
$\mathbb{E}[\left\lvert f\right\rvert^{q}\cdot\log\left\lvert f\right\rvert^{q}]={\int_{S_{3}}}\left\lvert f(x)\right\rvert^{q}\cdot\log(\left\lvert f(x)\right\rvert^{q})\pi(\mathrm{d}x)=-{\int_{S_{3}}}\rho(\left\lvert f(x)\right\rvert^{q})\pi(\mathrm{d}x)$
@@
$\mathbb{E}[\left\lvert f\right\rvert^{q}]\cdot\log\mathbb{E}[\left\lvert f\right\rvert^{q}]$ $=\left(\pi(S_{1})+\int_{S_{3}}\left\lvert f(x)\right\rvert^{q}\pi(\mathrm{d}x)\right)\cdot\log\left(\pi(S_{1})+\int_{S_{3}}\left\lvert f(x)\right\rvert^{q}\pi(\mathrm{d}x)\right)$
@@
$\mathbb{E}[\left\lvert f\right\rvert^{q}\cdot\log\left\lvert f\right\rvert^{q}]\geq-e^{-1}\cdot\pi(S_{3}).$ (10)
@@
$\mathbb{E}[\left\lvert f\right\rvert^{q}]\cdot\log\mathbb{E}[\left\lvert f\right\rvert^{q}]\geq\pi(S_{1})\cdot\log\left(\frac{1}{\pi(S_{1})}\right)-\pi(S_{3}).$ (11)
@@
$\pi(S_{3})$ $\geq\frac{\mathcal{D}(S_{1},S_{2})^{q}}{\mathtt{s}_{q}^{-1}+(1+e^{-1})\cdot\mathcal{D}(S_{1},S_{2})^{q}}\cdot\max\left\{\pi(S_{1})\cdot\log\left(\frac{1}{\pi(S_{1})}\right),\pi(S_{2})\cdot\log\left(\frac{1}{\pi(S_{2})}\right)\right\}$
$\geq\frac{\mathcal{D}(S_{1},S_{2})^{q}}{\mathtt{s}_{q}^{-1}+(1+e^{-1})\cdot\mathcal{D}(S_{1},S_{2})^{q}}\cdot\min\{\pi(S_{1}),\pi(S_{2})\}\cdot\log\left(\frac{1}{\min\{\pi(S_{1}),\pi(S_{2})\}}\right)$
$\geq\frac{\mathcal{D}(S_{1},S_{2})^{q}}{\mathtt{s}_{q}^{-1}+(1+e^{-1})\cdot\mathcal{D}(S_{1},S_{2})^{q}}\cdot\frac{\min\{\pi(S_{1}),\pi(S_{2})\}}{2}\cdot\log\left(\frac{2}{\min\{\pi(S_{1}),\pi(S_{2})\}}\right)$
@@
Notably, the $L^{q}$-Poincaré and $L^{q}$-log-Sobolev inequalities respectively lead to three-set isoperimetric inequalities with $\Upsilon(t)\sim\mathtt{p}_{q}t^{q}/2^{2q}$ and $\Upsilon(t)\sim\mathtt{s}_{q}t^{q}$ as $t\to 0$. Here, the sign $\sim$ means that the two quantities are equivalent as $t\to 0$, in the sense that their ratio converges to a finite constant. This regime is particularly relevant in high-dimensional settings, where $t$ typically scales inversely with the dimension, as per the discussion following Corollary 3.12. In contrast, when $\pi$ is log-concave, $\Upsilon(t)\sim t$ as $t\to 0$ (Andrieu et al., 2024, Subsection 3.2).
In many settings, one can only verify a close coupling condition on a high probability region $K\subsetneq\mathbb{R}^{d}$, which leads to lower bounds on the $s$-conductance (or restricted conductance) rather than the usual conductance; see, for example, Wu et al. (2022); Chen and Gatmiry (2023); Chewi et al. (2021); Dwivedi et al. (2019); Mangoubi and Vishnoi (2019); Mou et al. (2022a); Chen et al. (2020). In certain cases, obtaining such bounds requires working with a three-set isoperimetric inequality for the distribution $\pi$ truncated to $K$. For log-concave distributions, it is known that truncation preserves three-set isoperimetric inequalities; see Chen et al. (2020, Lemma 15) and Dwivedi et al. (2019, Proof of Lemma 6), both relying on Cousins and Vempala (2014, Theorem 4.4). Beyond the log-concave setting, however, it is unclear whether three-set isoperimetric inequalities are preserved under truncation. Extending the three-set isoperimetric inequalities associated with global Poincaré or log-Sobolev inequalities to truncated distributions appears delicate and is left for future work.
@@ 4.2 Close coupling conditions for the Gibbs kernels
In what follows, let $d\geq 2$. Here, we establish close coupling conditions for the systematic-scan and random-scan Gibbs kernels, assuming that the conditional distributions of the target distribution $\pi$ are TV continuous (as per Definition 2.5). The following results rely on sequential coupling arguments; see Roberts and Rosenthal (2013) for a formal construction of such couplings, and Marton (2004); Wang and Wu (2014) for related coupling techniques in Wasserstein distance. We begin with the systematic-scan Gibbs sampler.
@@
$\delta=\frac{1}{(Md)^{1/\beta}},\qquad\varepsilon=\frac{1}{4}.$
@@
By construction, $U\sim P_{\mathrm{SS}}(x,\cdot)$ and $V\sim P_{\mathrm{SS}}(y,\cdot)$. Moreover, observe that for all $k\in[d]$, $U^{\left(k\right)}=\left(U_{1:k},x_{k+1:d}\right)$ and $V^{\left(k\right)}=\left(V_{1:k},y_{k+1:d}\right)$. Using this, (12) can be rewritten as
@@
We now turn our attention towards obtaining a close coupling for the random-scan Gibbs sampler. However, one issue immediately becomes apparent: when $x,y\in\mathbb{R}^{d}$ do not share a coordinate (i.e., when xyx-y has a non-zero component in each coordinate direction), the supports of $P_{\mathrm{RS}}(x,\cdot)$ and $P_{\mathrm{RS}}(y,\cdot)$ are disjoint. In this case,
@@
To circumvent this issue, we adopt a similar approach to Narayanan and Srivastava (2022) by establishing a close-coupling condition for the iterated kernel $P_{\mathrm{RS}}^{N}$. Notably, when $N\geq d$, $P_{\mathrm{RS}}^{N}(x,\cdot)$ is globally supported for all $x\in\mathbb{R}^{d}$. It turns out that $N\sim d\cdot\log d$ as $d\to\infty$ is indeed the correct scaling in order to get a non-degenerate close coupling condition for the iterated random-scan Gibbs kernel, as formalized in the next lemma. Let $\lceil\cdot\rceil$ denote the ceiling function. The proof can be found in Appendix A.
@@
Let $I_{1},\dots,I_{N}\sim\mathrm{Unif}[d]$ be independent and identically distributed. Denote by $\Xi$ the set of unique values attained by $I_{1},\dots,I_{N}$. If $N=\lceil 2d\cdot\log d\rceil$, then
@@
Lemma 4.5 is closely related to the classic coupon collector problem in probability (Erdős and Rényi, 1961). Using the same notation, observe that when $N=d$,
@@
as $d\to\infty$. This follows from Stirling’s formula. The probability of observing all $d$ unique values in exactly $d$ draws thus decreases exponentially with the dimension $d$.
@@
Suppose that the conditional distributions of $\pi$ are $(M,\beta)$-TV continuous. Let $N_{d}=\lceil 2d\cdot\log d\rceil$. Then,
$\left\|P_{\mathrm{RS}}^{N_{d}}(x,\cdot)-P_{\mathrm{RS}}^{N_{d}}(y,\cdot)\right\|_{\textsc{TV}}\leq 1-\frac{1}{2}\left(1-M\left\lvert x-y\right\rvert^{\beta}\right)^{N_{d}}\qquad\forall x,y\in\mathbb{R}^{d}.$
@@
$\delta=\frac{1}{(MN_{d})^{1/\beta}},\qquad\varepsilon=\frac{1}{8}.$
@@
By construction, $U\sim P_{\mathrm{RS}}^{N_{d}}(x,\cdot)$ and $V\sim P_{\mathrm{RS}}^{N_{d}}(y,\cdot)$. As before, denote by $\Xi$ the set of unique values attained by $I_{1},\dots,I_{N_{d}}$. When $x$ and $y$ differ in at least two coordinates, the event $\left\{U=V\right\}$ has a non-zero probability of occurring only if $\Xi=[d]$, i.e., when all coordinates have been updated at least once by $P_{\mathrm{RS}}^{N_{d}}$. More generally, for all $x,y\in\mathbb{R}^{d}$,
@@
By Lemma 4.5, $\mathbb{P}(\Xi=\left[d\right])\geq 1/2$ when $N_{d}=\lceil 2d\cdot\log d\rceil$. Hence, for this value of $N_{d}$, we get
@@
$\leq M\left\lvert x-y\right\rvert^{\beta},$
where the subscript in the expectation indicates that it is taken with respect to the joint distribution of $U^{(k-1)},V^{(k-1)},I_{k}$. Recalling (15) and (16), we finally obtain
@@ 4.3 Mixing time bounds for the Gibbs sampler
By combining the three-set isoperimetric inequalities from Subsection 4.1 with the close coupling conditions from Subsection 4.2, we can bound the conductance of the systematic-scan and random-scan Gibbs kernels. These results apply to target distributions that satisfy an $L^{q}$-Poincaré or $L^{q}$-log-Sobolev inequality, and whose conditional distributions are TV continuous. Throughout this subsection, we use $\lesssim$ to denote inequalities that hold up to a universal constant factor.
@@
Consider a target distribution $\pi$ that satisfies $\mathrm{PI}_{q}$ with constant $\mathtt{p}_{q}$. Suppose that the conditional distributions of $\pi$ are $(M,\beta)$-TV continuous. Then, the conductance of the lazy systematic-scan Gibbs kernel is bounded by
$\Phi(\widetilde{P}_{\mathrm{SS}})\gtrsim\min\left\{1,\frac{\mathtt{p}_{q}}{2^{2q}(Md)^{q/\beta}}\right\}.$
@@
$\tau(\zeta,\widetilde{P}_{\mathrm{SS}})\lesssim\max\left\{1,\frac{2^{4q}(Md)^{2q/\beta}}{\mathtt{p}_{q}^{2}}\right\}\cdot\log\left(\frac{\sqrt{\Omega}}{\zeta}\right).$
@@
Combining Corollary 3.12, Proposition 4.1, and Proposition 4.4 yields a lower bound on the conductance of $P_{\mathrm{SS}}$. By (5), this directly implies the stated bound on the conductance of $\widetilde{P}_{\mathrm{SS}}$. Since $\widetilde{P}_{\mathrm{SS}}$ is lazy, the bound on the mixing time follows from Corollary 3.7. ∎
@@
Consider a target distribution $\pi$ that satisfies $\mathrm{PI}_{q}$ with constant $\mathtt{p}_{q}$. Suppose that the conditional distributions of $\pi$ are $(M,\beta)$-TV continuous. Let $N_{d}=\lceil 2d\cdot\log d\rceil$. Then, the conductance of the iterated random-scan Gibbs kernel is bounded by
$\Phi\left(P^{N_{d}}_{\mathrm{RS}}\right)\gtrsim\min\left\{1,\frac{\mathtt{p}_{q}}{2^{2q}(MN_{d})^{q/\beta}}\right\}.$
@@
$\lambda_{2}\left(P^{N_{d}}_{\mathrm{RS}}\right)\gtrsim\min\left\{1,\frac{\mathtt{p}_{q}^{2}}{2^{4q}(MN_{d})^{2q/\beta}}\right\}.$
@@
$\tau\left(\zeta,P^{N_{d}}_{\mathrm{RS}}\right)\lesssim\max\left\{1,\frac{2^{4q}(MN_{d})^{2q/\beta}}{\mathtt{p}_{q}^{2}}\right\}\cdot\log\left(\frac{\sqrt{\Omega}}{\zeta}\right).$
@@
Combining Corollary 3.12, Proposition 4.1, and Proposition 4.6 yields the lower bound on the conductance. The spectral gap is bounded using Theorem 3.8. Since $P_{\mathrm{RS}}$ is reversible and positive semi-definite, the same holds for $P^{N_{d}}_{\mathrm{RS}}$. The bound on the mixing time then follows from Corollary 3.7. ∎
We now focus on bounding the conductance and spectral gap of the random-scan Gibbs kernel $P_{\mathrm{RS}}$. Recall that $P_{\mathrm{RS}}$ is reversible and positive semi-definite. As noted by Narayanan and Srivastava (2022, Lemma 5.3), for any reversible Markov kernel $P$ and for all $N\geq 1$,
@@
Let $\sigma(P)$ denote the second largest eigenvalue of $P$, so that $\lambda_{2}(P)=1-\sigma(P)$. If $P$ is positive semi-definite, $\sigma\left(P^{N}\right)=\sigma(P)^{N}$, which implies that
@@
$\Phi(P_{\mathrm{RS}})\gtrsim\frac{1}{d\cdot\log d}\cdot\min\left\{1,\frac{\mathtt{p}_{q}}{2^{2q}(3Md\cdot\log d)^{q/\beta}}\right\}.$
@@
$\lambda_{2}(P_{\mathrm{RS}})\gtrsim\frac{1}{d\cdot\log d}\cdot\min\left\{1,\frac{\mathtt{p}_{q}^{2}}{2^{4q}(3Md\cdot\log d)^{2q/\beta}}\right\}.$
@@
$\tau(\zeta,P_{\mathrm{RS}})\lesssim d\cdot\log d\cdot\max\left\{1,\frac{2^{4q}(3Md\cdot\log d)^{2q/\beta}}{\mathtt{p}_{q}^{2}}\right\}\cdot\log\left(\frac{\sqrt{\Omega}}{\zeta}\right).$
These results demonstrate that the systematic-scan and random-scan Gibbs samplers exhibit polynomial mixing time dependence on the dimension $d$ for a large class of target distributions. Note that a single step of the systematic-scan kernel $P_{\mathrm{SS}}$ updates all $d$ coordinates, whereas the random-scan kernel $P_{\mathrm{RS}}$ only updates one. Even after accounting for this difference, a residual $\log d$ factor remains in the mixing time bounds for random-scan. This additional $\log d$ factor highlights a modest penalty compared to systematic scan for well-behaved target distributions. This aligns with prior heuristic observations about the efficiency of these sampling schemes (Diaconis, 2013) and arises naturally from spectral bounds on mixing times (Gaitonde and Mossel, 2026).
By Proposition 4.3, all results in this subsection extend to distributions satisfying $\mathrm{LSI}_{q}$ instead of $\mathrm{PI}_{q}$, with the conductance bounds remaining equivalent up to constant factors. Consequently, the bounds on the mixing time in total variation distance remain unchanged when replacing a Poincaré inequality with a log-Sobolev inequality of the same order.
However, improvements may be possible when using a different norm to define the mixing time. One possible approach is to adapt spectral and conductance profile methods under an $L^{q}$-log-Sobolev inequality, following Andrieu et al. (2024); Chen et al. (2020). In $\chi^{2}$-distance, these methods exploit the logarithmic factor in the function $\Psi$ from Proposition 4.3 to improve the dependence of the mixing time on the warm start parameter, reducing it from $\log$ to $\log\log$ (see Corollary 3.7). At a high level, this would require sharpening Corollary 3.12 along the lines of Andrieu et al. (2024, Corollary 16) to obtain tighter control of the conductance profile, and then plugging this refinement into the mixing time bounds of Andrieu et al. (2024, Theorem 8). Implementing this approach would require balancing several interacting quantities, such as the order of the Poincaré or log-Sobolev inequality, the form of the function $\Psi$ in the associated three-set isoperimetric inequality, and the exponent in the regularity assumption for the conditional distributions (Definition 2.5), resulting in less transparent expressions. Moreover, spectral profile methods rely on reversibility and do not improve upon standard spectral gap bounds for distributions that satisfy only a Poincaré inequality (Andrieu et al., 2024, Subsection 3.3). For these reasons, we do not pursue this direction further here.